OpenStudy (anonymous):

HELP 7. An engineer is surveying a building with a device that is 6 feet above the ground. He is 60 feet from the building and is looking with an angle of elevation of 45°. How tall is the building?

OpenStudy (anonymous):

not exactly clear to me what this means, but i would say 60 feet plus another 6 feet for a total of 66

OpenStudy (anonymous):

I'm reading... trying to reconcile your approach with satellite's...

OpenStudy (anonymous):



OpenStudy (anonymous):

why did you type: tan(60)(45) = 77.94 ? You lost me on that step.

OpenStudy (anonymous):

I see the error (finally). You should have found 60 * tan(45)... you did it like tan(60)(45).... so you found the tangent of a 60 degree angle, not a 45 degree angle.

OpenStudy (anonymous):

opp / adj = tan(45) x / 60 = tan(45) x = 60 * tan(45) x = 60 * 1 x = 60 But don't forget to add the extra 6 feet, since the way x is measured, it is only from the top of the survey device.

OpenStudy (anonymous):

So as satellite73 showed, it would be 60 + 6 = 66 feet. Take a look at that other problem you attached... it was confusing, maybe, since both the angle and the distance were 50. But you take the tangent of the angle, and that is equal to "opp / adj". So in this currrent problem, you have to take tan(45) and set that equal to opp/adj.

OpenStudy (anonymous):

@charlotteakina, does that make sense? Do you see where you had the error?

OpenStudy (anonymous):

yes, sorry, I stepped away. that's correct... glad you see the concept also :) Good luck!